OpenStudy (anonymous):

Find the solution set of the equation by using substitution to transform the equation into a quadratic equation that you can solve. w - sqrtw - 6 = 0

Directrix (directrix):

Try this: Let w = x^2 Then this --> w - sqrtw - 6 = 0 becomes x^2 - x - 6 = 0 Solve that and post what you get. There will be a little more work to do after that. @shoplovepc